# Deep Learning of Force Manifolds from the Simulated Physics of Robotic Paper Folding

### 05 Feb 2024 v3

@@ Abstract
Robotic manipulation of slender objects is challenging, especially when the induced deformations are large and nonlinear. Traditionally, learning-based control approaches, such as imitation learning, have been used to address deformable material manipulation. These approaches lack generality and often suffer critical failure from a simple switch of material, geometric, and/or environmental (e.g., friction) properties. This article tackles a fundamental but difficult deformable manipulation task: forming a predefined fold in paper with only a single manipulator. A data-driven framework combining physically-accurate simulation and machine learning is used to train a deep neural network capable of predicting the external forces induced on the manipulated paper given a grasp position. We frame the problem using scaling analysis, resulting in a control framework robust against material and geometric changes. Path planning is then carried out over the generated “neural force manifold” to produce robot manipulation trajectories optimized to prevent sliding, with offline trajectory generation finishing 15$\times$ faster than previous physics-based folding methods. The inference speed of the trained model enables the incorporation of real-time visual feedback to achieve closed-loop sensorimotor control. Real-world experiments demonstrate that our framework can greatly improve robotic manipulation performance compared to state-of-the-art folding strategies, even when manipulating paper objects of various materials and shapes.
@@
This article is motivated by the need for efficient robotic folding strategies for stiff materials such as paper. Previous robot folding strategies have focused primarily on soft materials (e.g., cloth) possessing minimal bending resistance or relied on multiple complex manipulators and sensors, significantly increasing computational and monetary costs. In contrast, we formulate a robust, data-driven, physics-based method capable of folding papers of varying stiffness with a single manipulator. The proposed folding scheme is limited to papers of homogeneous material and folding along symmetric centerlines. Future work will involve formulating efficient methods for folding along arbitrary geometries and preexisting creases.
@@
robotic deformable manipulation, deep neural networks, data-driven models, closed-loop sensorimotor control
@@ I Introduction
Prior research has focused primarily on manipulating either cloth [1, 2, 3, 4, 5, 6] or ropes [7, 8, 9, 10, 11, 12, 13, 14] and, as a result, the challenge of robotically manipulating many other deformable objects still lacks robust solutions. This article addresses a particularly difficult deformable manipulation task — folding paper. Paper is similar to cloth but typically possesses a significantly higher bending stiffness and a slippery surface. Therefore, when compared to folding garments and fabrics, the folding of paper requires more delicate and insightful manipulations. In fact, in our experiments we observe that state-of-the-art methods for robotic fabric/cloth folding [15, 16, 17] perform poorly when transferred to paper.
To tackle these challenges, we propose a framework that combines physically accurate simulation, scaling analysis, and machine learning to generate folding trajectories optimized to prevent sliding. With scaling analysis, we make the problem non-dimensional, resulting in both dimensionality reduction and generality. This allows us to train a single nondimensionalized neural network, whose outputs are referred to as a neural force manifold (NFM), to continuously approximate a scaled force manifold sampled purely from simulation. Compared to numerical models that require the entire geometric configuration of the paper, NFMs map the external forces of the paper given only the grasp position. Therefore, we can generate trajectories optimized to minimize forces (and thus minimize sliding) by applying path planning algorithms. Furthermore, the nondimensionality of the NFM allows us to generate trajectories for paper of various materials and geometric properties even if such parameters were not present in the training dataset. We show that our approach is capable of folding paper on extremely slick surfaces with little-to-no sliding (Fig. 1(b)).
@@
    we formulate a solution for folding homogeneous material papers along symmetrical centerlines in a physically robust manner using scaling analysis, resulting in complete generality concerning the modulus and density of the material, size of the paper, and environmental properties (e.g., friction);
@@
    we train a neural network with non-dimensional simulation data forming a fast and accurate model that can generate a descriptive force manifold for trajectory optimization;
@@
    we exploit the high inference speed of our trained model with a perception system to construct a robust and efficient closed-loop sensorimotor control algorithm for the folding task in near real-time, and finally
  4. 4.
    we demonstrate full sim2real realization through an extensive robotic case study featuring 280 experiments across paper sheets of various materials and shapes. We compare our method against both natural paper folding strategies as well as the previous state of the art in robotic rectangular fabric folding [17, 15].
@@
The remainder of the article is organized as follows: We begin with a review of related work in Sec. II. A brief description of the folding problem is presented in Sec. III. The formulation of a reduced-order physics-based model is discussed in Sec. IV, where we formulate the folding problem using scaling analysis. In Sec. V, we formulate our learning framework as well as algorithms for optimal path planning. Next, in Sec. VI, we introduce our robotic system and formulate our closed-loop visual feedback pipeline. Experimental results for a robot case study and analysis of the results are given in Sec. VII. Finally, we make concluding remarks and discuss the potential of future research avenues in Sec. VIII.
@@ II Related Work
Vision-based approaches involve folding deformable materials by generating folding motions purely from visual input. These approaches are usually common for folding clothes [4, 20, 2] as they are extremely soft, which results in the easy predictability of their deformation state given a particular action. Such approaches can be effective and rather simple to implement, but do not transfer well to paper folding as paper possesses a much higher stiffness when compared to fabric and will attempt to restore its natural, undeformed state if not properly handled.
@@
Model-based approaches, where the model can either be known or learned, often use model predictive control to manipulate the deformable object. Learned models involve learning the natural dynamics of deformable objects through random perturbations [28]. These models are generally fast, but they can be inaccurate when experiencing new states. Theoretical models are often formulated to be as physically accurate as possible, which enables the direct application of their predictive power in the real world. Examples of this have been published for both strip folding [15, 17] and garment folding [29]. Physical models are often constructed using energy-based formulations [30, 31, 32], where various elastic energies are computed based on the topological properties of the simulated objects to solve their deformed shape under manipulation. For example, Wakamatsu and Hirai [33] modeled deformable linear objects (rods) with flexure (bending), torsion, and extension (stretching), while Jia et al. [34] introduced manipulation as a potential energy to compute the deformations of deformable planar objects. However, theoretical models are usually quite expensive to run and must often face a trade-off between accuracy and efficiency.
@@ III Problem Statement
Figure 2: Folding sheets of paper. The manipulation process involves (a) the initial state, where the paper lies flat on the substrate, followed by (b) the folding state, where the manipulated end is moved to the “crease target” line $C$, and finally (c) the folded state, which involves forming the desired crease on the paper.
@@ IV-A Reduced-Order Model Representation
Paper is a unique deformable object. Unlike cloth, its surface is developable [38]; i.e., the surface can bend but not stretch. Furthermore, shear deformations are not of particular importance as paper possesses a negligible thickness to length ratio. Therefore, the primary nonlinear deformation when folding paper in our scenario is bending deformation. We postulate that the nonlinear behaviors of paper arise primarily from a balance of bending and gravitational energies: $\epsilon_{b}\sim\epsilon_{g}$.
@@
where $\rho$ is its volume density and $H$ is its vertical height above the rigid substrate.
@@ IV-B Discrete Differential Geometry Numerical Model
Subsequent to the the pioneering work on physics-based modeling and simulation of deformable curves, surfaces, and solids in computer graphics [39, 40, 41], the community has shown impressive results using DDG-based simulation frameworks. For example, the Discrete Elastic Rod (DER) [37] framework has shown efficient and physically accurate simulation of deformable linear objects in various scenarios including knot tying [42], helix bifurcations [43], coiling of rods [44], and flagella buckling [45]. Given this success, we use the DER to model the centerline of the paper as a 2D planar rod undergoing bending deformations.
As shown in Fig. 3(c), the discrete model is comprised of $N+1$ nodes, $\mathbf{q}_{i}$ ($0\leq i\leq N$). Each node $\mathbf{q}_{i}$ represents two degrees of freedom (DOF): position along the $x$ and the $z$ axes. This results in a $22N+$ 2-sized DOF column vector $\mathbf{q}=[\mathbf{q}_{0},\mathbf{q}_{1},...,\mathbf{q}_{N}]^{T}$ representing the configuration of the paper sheet. Initially, all the nodes of the paper are located in a line along the $x$-axis in the paper’s undeformed state. As the robotic manipulator imposes boundary conditions on the end node $\mathbf{q}_{N}$, portions of the paper deform against the substrate as shown in Fig. 4(a). We compute the DOFs as a function of time $\mathbf{q}(t)$ by integrating the equations of motion at each DOF.
@@
where $k_{b}=\frac{Ewh^{3}}{12}$ is the bending stiffness, $w$ and $h$ are the width and thickness, respectively, $\phi_{i}$ is the “turning angle” at a node as shown in Fig. 3(d), and $\phi^{0}_{i}$ is the undeformed turning angle (00 for paper). Finally, since deformations are limited to a 2D plane, we can ignore twisting energies. The total elastic energy is therefore $E_{el}=E_{s}+E_{b}$.
@@ IV-C Generalized Solution and Scaling Analysis
We first denote several quantities to describe the deformed configuration of the paper. We introduce a point $\mathbf{q}_{C}$, which is the node that connects the suspended ($z>0$) and contact regions ($z=0$) of the paper. We focus solely on the suspended region as deformations occur primarily in this region. An origin $\mathbf{o}$ is defined for our 2D plane which is located at the initial manipulated end $\mathbf{q}_{N}$ as shown in Fig. 4(a). For the manipulated end, the robot end-effector imposes a position $\mathbf{q}_{N}=(x,z)$ and an orientation angle $\alpha$ to control the pose of the manipulated end as shown in Fig. 4(a). We impose a constraint that the curvature at the manipulated end is always zero so that sharp bending deformations are prevented, which is crucial to preventing permanent deformations during the folding process. On the connective node $\mathbf{q}_{C}$, the tangent is always along the $x$-axis. With these definitions, we can now modify (7) with the following constraints:
@@
Figure 4: (a) Side view of a symmetrical paper during folding with coordinate frame and relevant notations. (b) Sampled $\lambda$ forces for a particular $\bar{l}_{s}$ of 4.10. This showcases one of the sampled “partial” force manifolds that we use to train our neural network on.
@@
To proceed, we reduce the dimensionality of the problem by applying scaling analysis. According to the Buckingham $\pi$ theorem, we construct five unitless groups: $\bar{x}=x/L_{gb}$; $\bar{z}=z/L_{gb}$; $\bar{l}_{s}=l_{s}/L_{gb}$; $\alpha$; and $\lambda=F_{t}/F_{n}$, where $L_{gb}$ is the gravito-bending length (4). This results in the following unitless formulation of (9):
@@
Note that the mapping $\mathcal{F}$ is now independent of quantities with units; e.g., material and geometric properties of the paper. As the dimensionality of our problem has been reduced significantly, we can now express $\lambda$ as a function of just two parameters $\bar{x},\bar{z}$. Therefore, training a neural network to model $\mathcal{F}$ is now trivial as non-dimensionalized simulation data from a single type of paper can be used. Furthermore, the low dimensionality of $\mathcal{F}$ allows us easily to visualize the $\lambda$ landscape along a non-dimensional 2D-plane.
@@ V-C Constructing the Neural Force Manifold
Figure 6: Example of our perception system with a top down view of the folding procedure. (a) Shows the intuitive baseline results while (b) shows our open-loop algorithm for $L_{gb}=0.048$ and $C=0.25$ m. As in Fig. 2, the solid green line indicates the desired end effector position while the dashed blue line indicates the crease location. For this case, we observe that the intuitive baseline suffers from considerable sliding while our open-loop algorithm has near-perfect performance.
@@ 13      $\tau^{*}\leftarrow$ perform trajectory smoothing on $\tau^{*}$ return $\tau^{*}$
V-D Path Planning over the Neural Force Manifold
Given the discretized manifold $\mathcal{M}$, we can now generate optimal trajectories through traditional path planning algorithms. Indeed, we can find that there exist two local minima regions (dark blue in Fig. 5(a)) in the neural force manifold $\mathcal{M}$. However, note that these two minima regions are not connected, which means improper local optimization may result in undesired traversal through high force regions later. As mentioned previously, prior mechanics-based work for folding shell-like structures (cloth) have used either physical simulations or energy-based optimization to compute the optimal subsequent grasp based solely on the current status of the manipulated object [15, 17]. We show that this local optimization approach performs poorly for paper folding in Sec. VII. In contrast, we generate globally optimized trajectories that take into account both current and future states of the paper. To do so, we define an optimal trajectory $\tau^{*}$ as one that reaches the goal state while minimizing the sum of $\lambda$:
$\tau^{*}=\operatorname*{arg\,min}_{\tau\in\mathcal{T}}\sum^{i=L-1}_{i=0}\lambda_{i},$ (14)
Input: $(x_{s},z_{s}),(x_{g},z_{g}),L_{gb},\delta,N,\mathcal{F}_{\textrm{NN}}$
crease paper with roller
Algorithm 2 Closed-loop Control Pseudocode

### 20 Feb 2024 v4

# Learning Neural Force Manifolds for Sim2Real Robotic Symmetrical Paper Folding

@@ Abstract
Robotic manipulation of slender objects is challenging, especially when the induced deformations are large and nonlinear. Traditionally, learning-based control approaches, such as imitation learning, have been used to address deformable material manipulation. These approaches lack generality and often suffer critical failure from a simple switch of material, geometric, and/or environmental (e.g., friction) properties. This article tackles a fundamental but difficult deformable manipulation task: forming a predefined fold in paper with only a single manipulator. A sim2real framework combining physically-accurate simulation and machine learning is used to train a deep neural network capable of predicting the external forces induced on the manipulated paper given a grasp position. We frame the problem using scaling analysis, resulting in a control framework robust against material and geometric changes. Path planning is then carried out over the generated “neural force manifold” to produce robot manipulation trajectories optimized to prevent sliding, with offline trajectory generation finishing 15$\times$ faster than previous physics-based folding methods. The inference speed of the trained model enables the incorporation of real-time visual feedback to achieve closed-loop model-predictive control. Real-world experiments demonstrate that our framework can greatly improve robotic manipulation performance compared to state-of-the-art folding strategies, even when manipulating paper objects of various materials and shapes.
@@
This article is motivated by the need for efficient robotic folding strategies for stiff materials such as paper. Previous robot folding strategies have focused primarily on soft materials (e.g., cloth) possessing minimal bending resistance or relied on multiple complex manipulators and sensors, significantly increasing computational and monetary costs. In contrast, we formulate a robust, sim2real, physics-based method capable of folding papers of varying stiffness with a single manipulator. The proposed folding scheme is limited to papers of homogeneous material and folding along symmetric centerlines. Future work will involve formulating efficient methods for folding along arbitrary geometries and preexisting creases.
@@
deformable object manipulation, sim2real paper folding, data-driven models, closed-loop model-predictive control
@@ I Introduction
Prior research has focused primarily on manipulating either cloth or ropes [1, 2], and, as a result, the challenge of robotically manipulating many other deformable objects still lacks robust solutions. This article addresses a particularly difficult deformable manipulation task — folding paper. Paper is similar to cloth but typically possesses a significantly higher bending stiffness and a slippery surface. Therefore, when compared to folding garments and fabrics, the folding of paper requires more delicate and insightful manipulations. In fact, in our experiments, we observe that state-of-the-art methods for robotic fabric/cloth folding [3, 4, 5] perform poorly when transferred to paper.
To tackle these challenges, we propose a framework that combines physically accurate simulation, scaling analysis, and machine learning to generate folding trajectories optimized to prevent sliding. With scaling analysis, we make the problem non-dimensional, resulting in both dimensionality reduction and generality. This allows us to train a single nondimensionalized neural network, whose outputs are referred to as a neural force manifold (NFM), to continuously approximate a scaled force manifold sampled purely from simulation. Compared to numerical models that require the entire geometric configuration of the paper, NFMs map the external forces of the paper given only the grasp position. Therefore, we can generate trajectories optimized to minimize forces (and thus minimize sliding) by applying path-planning algorithms. Furthermore, the nondimensionality of the NFM allows us to generate trajectories for papers of various materials and geometric properties even if such parameters were not present in the training dataset. We show that our approach is capable of folding paper on extremely slick surfaces with little-to-no sliding (Fig. 1(b)).
@@
    We formulate a solution for folding materially homogeneous sheets of paper along symmetrical centerlines in a physically robust manner using scaling analysis, resulting in complete generality concerning the modulus and density of the material, size of the paper, and environmental properties (e.g., friction).
@@
    Next, we generate accurate non-dimensional simulation data to train a “neural force manifold” for optimal trajectory generation. We exploit the high inference speed of our trained model with a perception system to construct a robust and efficient closed-loop model-predictive control algorithm for the folding task in near real-time.
@@
    Finally, we demonstrate full sim2real realization through an extensive robotic case study featuring 360+ folding experiments involving paper sheets of various materials and shapes. We compare our method against both natural paper folding strategies as well as the previous state of the art in robotic rectangular fabric folding [5, 3].
@@
The remainder of the article is organized as follows: We begin with a review of related work in Sec. II. A brief description of the folding problem is presented in Sec. III. The formulation of a reduced-order physics-based model is discussed in Sec. IV, where we formulate the folding problem using scaling analysis. In Sec. V, we formulate our learning framework as well as algorithms for optimal path planning. Next, in Sec. VI, we introduce our robotic system and formulate our closed-loop visual feedback pipeline. Experimental results for a robot case study and analysis of the results are given in Sec. VII. Subsequently, Sec. VIII provides additional discussion regarding performing multiple folds and the importance of single-manipulator folding. Finally, we make concluding remarks and discuss the potential of future research avenues in Sec. IX.
@@ II Related Work
Vision-based approaches involve folding deformable materials by generating folding motions purely from visual input. These techniques are commonly applied to tasks such as folding clothes, where the primary focus is on detecting garment shape or key grasp points. Techniques for key feature extraction involve random decision trees [8], RGB-D sensing data analysis [9, 10], and fitting strategies where the detected state of deformed clothes are compared against precomputed shapes [11, 12]. Given the soft nature of clothes, subsequent manipulations are often formulated intuitively. While some prior research employs models to predict optimal manipulation sequences, these models are typically oversimplified and lack physical details [13]. Such approaches can be effective and rather simple to implement, but do not transfer well to paper folding as paper has a much higher stiffness than fabric and will attempt to restore its natural, undeformed state if not properly handled.
@@
Model-based approaches, where the model can either be known or learned, often use model-predictive control to manipulate the deformable object. Learned models involve learning the natural dynamics of deformable objects through random perturbations [21]. These models are generally fast, but they can be inaccurate when experiencing new states. Theoretical models are often formulated to be as physically accurate as possible, which enables the direct application of their predictive power in the real world. Examples of this have been published for both strip folding [3, 5] and garment folding [22]. Physical models are often constructed using energy-based formulations [1, 2, 23], where various elastic energies are computed based on the topological properties of the simulated objects to solve their deformed shape under manipulation. For example, Wakamatsu and Hirai [24] modeled deformable linear objects (rods) with flexure (bending), torsion, and extension (stretching), while Jia et al. [25] introduced manipulation as a potential energy to compute the deformations of deformable planar objects. However, theoretical models are usually quite expensive to run and must often face a trade-off between accuracy and efficiency.
@@ III Problem Statement
Figure 2: States of the paper during the folding process. The manipulation process involves two steps. The first (folding) step transitions the paper from the initial state (a), where the paper lies flat on the substrate, to the folding state (b), where the manipulated end is moved to the “crease target” line $C$. The second (creasing) step then transitions the paper from state (b) to the final folded state (c), which involves forming the desired crease on the paper.
@@ IV-A Reduced-Order Model Representation
Paper is a unique deformable object. Unlike cloth, its surface is developable [29]; i.e., the surface can bend but not stretch. Furthermore, shear deformations are not of particular importance as paper possesses a negligible thickness-to-length ratio. Therefore, the primary nonlinear deformation when folding paper in our scenario is bending deformation. We postulate that the nonlinear behaviors of paper arise primarily from a balance of bending and gravitational energies: $\epsilon_{b}\sim\epsilon_{g}$.
@@
where $\rho$ is its volume density, and $H$ is its vertical height above the rigid substrate.
@@ IV-B Discrete Differential Geometry Numerical Model
Subsequent to the pioneering work on physics-based modeling and simulation of deformable curves, surfaces, and solids in computer graphics [30, 31, 32], the community has shown impressive results using DDG-based simulation frameworks. For example, the Discrete Elastic Rods (DER) [28] framework has shown remarkable efficiency and physically validated accuracy when simulating deformable linear objects for a wide variety of scenarios including elastic coiling [33], deployment [34], helix bifurcation [35], overhand knot tightening [36, 37], overhand knot buckling [38], flagella buckling [39, 37], and dynamic cantilever beams [40]. Given this success, we too choose to use DER to model the centerline of the paper as a 2D planar rod undergoing bending deformations.
As shown in Fig. 3(c), the discrete model is comprised of $N+1$ nodes, $\mathbf{q}_{i}$ ($0\leq i\leq N$). Each node $\mathbf{q}_{i}$ represents two degrees of freedom (DOF): position along the $x$ and the $z$ axes. This results in a $22N+$ 2-sized DOF column vector $\mathbf{q}=[\mathbf{q}_{0},\mathbf{q}_{1},...,\mathbf{q}_{N}]^{T}$ representing the configuration of the paper sheet. Initially, all the nodes of the paper are located in a line along the $x$-axis in the paper’s undeformed state. As the robotic manipulator imposes boundary conditions on the end node $\mathbf{q}_{N}$, portions of the paper deform against the substrate, as shown in Fig. 4(a). We compute the DOFs as a function of time $\mathbf{q}(t)$ by integrating the equations of motion at each DOF.
@@
where $k_{b}=\frac{Ewh^{3}}{12}$ is the bending stiffness, $w$ and $h$ are the width and thickness, respectively, $\phi_{i}$ is the “turning angle” at a node (Fig. 3(d)), and $\phi^{0}_{i}$ is the undeformed turning angle (00 for paper). Finally, since deformations are limited to a 2D plane, we can ignore twisting energies. The total elastic energy is therefore $E_{el}=E_{s}+E_{b}$.
@@ IV-C Generalized Solution and Scaling Analysis
We first denote several quantities to describe the deformed configuration of the paper. We introduce a point $\mathbf{q}_{C}$, which is the node that connects the suspended ($z>0$) and contact regions ($z=0$) of the paper. We focus solely on the suspended region as deformations occur primarily in this region. An origin $\mathbf{o}$ is defined for our 2D plane which is located at the initial manipulated end $\mathbf{q}_{N}$, as shown in Fig. 4(a). For the manipulated end, the robot end-effector imposes a position $\mathbf{q}_{N}=(x,z)$ and an orientation angle $\alpha$ to control the pose of the manipulated end (Fig. 4(a)). We impose a constraint that the curvature at the manipulated end is always zero so that sharp bending deformations are prevented, which is crucial to preventing permanent deformations during the folding process. On the connective node $\mathbf{q}_{C}$, the tangent is always along the $x$-axis. With these definitions, we can now modify (7) with the following constraints:
@@
Figure 4: (a) Side view of a symmetrical paper during folding with coordinate frame and relevant notations. (b) Sampled $\lambda$ forces for a particular $\bar{l}_{s}$ of 4.10. This shows one of the sampled “partial” force manifolds that we use to train our neural network on.
@@
To proceed, we reduce the dimensionality of the problem by applying scaling analysis. According to the Buckingham $\pi$ theorem, we construct five unitless groups: $\bar{x}=x/L_{gb}$; $\bar{z}=z/L_{gb}$; $\bar{l}_{s}=l_{s}/L_{gb}$; $\alpha$, and $\lambda=F_{t}/F_{n}$, where $L_{gb}$ is the gravito-bending length (4). This results in the following unitless formulation of (9):
@@
Note that the mapping $\mathcal{F}$ is now independent of quantities with units; e.g., material and geometric properties of the paper. As the dimensionality of our problem has been reduced significantly, we can now express $\lambda$ as a function of just two parameters $\bar{x},\bar{z}$. Therefore, training a neural network to model $\mathcal{F}$ is now trivial as non-dimensionalized simulation data from a single type of paper can be used. Furthermore, the low dimensionality of $\mathcal{F}$ allows us to easily visualize the $\lambda$ landscape along a non-dimensional 2D-plane.
@@ V-C Constructing the Neural Force Manifold
V-D Path Planning over the Neural Force Manifold
Given the discretized manifold $\mathcal{M}$, we can now generate optimal trajectories through traditional path planning algorithms. Indeed, we find that there exist two local minima regions (dark blue in Fig. 5(a)) in the neural force manifold $\mathcal{M}$. However, note that these two minima regions are not connected, which means that improper local optimization may result in undesired traversal through high force regions later. As mentioned previously, prior mechanics-based efforts on folding shell-like structures (cloth) have used either physical simulations or energy-based optimization to compute the optimal subsequent grasp based solely on the current status of the manipulated object [3, 5]. We show that this local optimization approach performs poorly for paper folding in Sec. VII. By contrast, we generate globally optimized trajectories that take into account both current and future states of the paper. To do so, we define an optimal trajectory $\tau^{*}$ as one that reaches the goal state while minimizing the sum of $\lambda$:
$\tau^{*}=\operatorname*{arg\,min}_{\tau\in\mathcal{T}}\sum^{L-1}_{i=0}\lambda_{i},$ (14)